OpenStudy (anonymous):

is 2/3 a complex number?

OpenStudy (anonymous):

Is this a test?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Number

OpenStudy (anonymous):

Are there any imaginary numbers, then........

myininaya (myininaya):

its complex

myininaya (myininaya):

any number is complex

OpenStudy (anonymous):

2/3 +0*i

OpenStudy (anonymous):

any number is complex?

jimthompson5910 (jim_thompson5910):

Depends on which math class you are in. For elementary algebra classes, this is a real number (since complex is yet to be introduced). The number is complex in this context if it has an imaginary part to it. Note: overall, this number is complex since it's of the form a+bi where a=2/3 and b=0, but I get the feeling that the teacher is aiming for the first definition

OpenStudy (zarkon):

it's complex

jimthompson5910 (jim_thompson5910):

To be honest, the terminology is a bit strange when you encounter it. So teachers should address this issue a little better.

OpenStudy (zarkon):

it is not imaginary ... but is complex

myininaya (myininaya):

any number is complex it is in the form a+bi where a and b are real so 2/3=2/3+0*i 0 is a real number so 2/3 is like totally complex

OpenStudy (anonymous):

Let's just settle for rational....

myininaya (myininaya):

there should not be several definitions for complex all sets of numbers are subsets of the complex

jimthompson5910 (jim_thompson5910):

But when you don't know about complex numbers, everything is real (or it's not real). If something is "real" in an elementary algebra class, then it's just any old number But if some number is not real, then it's considered to be complex ex: the number 2 is real, but 2+i is complex. In an elementary algebra class, 2 would NOT be considered complex (even though it really is)

OpenStudy (zarkon):

it should be considered complex since it is

jimthompson5910 (jim_thompson5910):

i guess what they mean is the best description

myininaya (myininaya):

but if they don't know anything about complex numbers, then why is the teacher asking about complex numbers?

jimthompson5910 (jim_thompson5910):

my guess is that they were introduced to them halfway through

myininaya (myininaya):

thats a horrible thing to do it sounds like a way to confuse the students

jimthompson5910 (jim_thompson5910):

tell me about it, this is probably why many students hate algebra

OpenStudy (anonymous):

algebra rocks!

OpenStudy (zarkon):

they are specifically asking if it is a complex number...the only answer is yes

myininaya (myininaya):

i totally agree with zarkon but it seems wolfram is not on our side

OpenStudy (anonymous):

Could be the op is asking on their own account...

jimthompson5910 (jim_thompson5910):

but I've seen plenty of problems ask people to determine the nature of the roots For instance, the equation x^2+4x+4=0 clearly has 2 real roots. But would you say that it also has complex roots? Sure the root itself (2) is complex, but an elementary algebra teacher would say that this root is real (a real double root actually)

jimthompson5910 (jim_thompson5910):

and yes even wolfram even says it's not complex, so it's probably thinking along the same lines of reasoning as elementary algebra teachers

OpenStudy (anonymous):

Negative then complex just came from a desire to have answers for every polynomial

OpenStudy (zarkon):

\[\mathbb{R}\subset \mathbb{C}\] every real number is complex

myininaya (myininaya):

i don't get why wolfram says its not complex. i think they say if you have b^2-4ac>0 gives you two real roots instead of saying two complex is because they are trying to narrow your set of numbers down

jimthompson5910 (jim_thompson5910):

it's looking for the imaginary part, since it's not finding it, it's saying it's not complex

OpenStudy (anonymous):

I suppose there is no reason to call it complex unless there is a reason to...

OpenStudy (zarkon):

wolfram is incorrect in this case

myininaya (myininaya):

i don't like that

myininaya (myininaya):

wolfram is evil and cannot be trusted

jimthompson5910 (jim_thompson5910):

lol

jimthompson5910 (jim_thompson5910):

even though it's used constantly

OpenStudy (anonymous):

How to confuse Wolfram http://www.wolframalpha.com/input/?i=plot+x%3D2%2F3%2C+x%3D+2%2F3+%2B+0i

OpenStudy (anonymous):

Links are busted..

OpenStudy (zarkon):

From wolfram:"The complex numbers are the field C of numbers of the form x+iy, \ where x and y are real numbers and i is the imaginary unit equal to \ the square root of -1, sqrt(-1)." looks like they are not using their own definition

OpenStudy (anonymous):

Next time they ask me for an interval I'm gonna put C

OpenStudy (anonymous):

Who needs R, lol